solve the following systems of linear equations using any algebraic method.check work if possible
8x-6y=12
12x-9y=18

Answer by macston(5194)

'
8x-6y=12
8x=6y+12
4x=3y+6
.
12x-9y=18
12x=9y+18
4x=3y+6 The equations are the same.
4x-6=3y
(4/3)x-2=y
The solution is all points on the line y=(4/3)x-2
.